# Compounding Vulnerability: Hub Removal Triggers Cascade Phase Transitions While Degrading Percolation Robustness in Scale-Free Networks

### 05 Mar 2026 v1

@@ Abstract
Hub removal in scale-free networks is known to degrade percolation robustness by raising the bond percolation threshold. We show that this same intervention simultaneously triggers a cascade phase transition under the Watts threshold model. In Barabási–Albert networks (N=2, [^0] N=2{,}000, $m=2$), removing the top 10% of nodes by degree raises $p_{c}$ from 0.34 to 0.90 ($168\%$), reducing robustness to random edge failure. Simultaneously, at cascade threshold $\varphi=0.22$, mean cascade size increases from $0.29\%$ to $20.6\%$ (200 trials), crossing from the subcritical to supercritical regime—compounding, rather than trading off, the percolation degradation.
Using a controlled experiment that independently varies hub presence and hub activation threshold on the same network topology, we demonstrate that hub-mediated cascade suppression is primarily dynamical: making hubs vulnerable without removing them produces 95% cascades versus 19% with removal and 0.28% at baseline. An intermediate condition (hub threshold $\varphi_{h}=0.05$) yields 85% cascades, revealing a continuous firewall gradient.
We operationalize Watts’ (2002) stable-node mechanism as a quantitative condition $k>1/\varphi$ and derive, under the configuration-model approximation, a closed-form expression for the post-removal cascade branching factor $z_{1}(\varphi,\alpha,m)$. This predicts a newly opened interval where the pre-removal network is subcritical but the post-removal network is supercritical. We define $\varphi^{*}$ as the upper supercritical boundary, $\varphi^{*}=\sup\{\varphi:z_{1}(\varphi)\geq 1\}$. The $z_{1}$ derivation confirms that at $\varphi=0.22$, the pre-removal network is subcritical ($z_{1}=0.850$) while the post-removal network is supercritical ($z_{1}=1.195$). Our results establish that hub manipulation creates compounding vulnerability: both percolation and cascade metrics worsen simultaneously, with no Pareto improvement achievable through uniform hub deletion on static, unweighted networks. These findings persist across N=500N=50010,00010{,}000 and, within the BA model family, require degree heterogeneity $\kappa>{\sim}10$; the effect vanishes in homogeneous networks (ER, WS) and its dependence on clustering remains an open question.
@@ I Introduction
Network resilience—the ability of a system to maintain function under stress—is a central problem of modern complexity science [1, 2, 3]. Yet “resilience” is routinely reported as if it were a scalar attribute of a network, an intrinsic capacity to withstand shocks. This scalarization is conceptually problematic: resilience is not a property of a graph $G$ alone, but of the pair $(G,S)$ consisting of a network and a specified stress model [4]. Conflating stress models produces what we term the resilience illusion—a category error that shapes engineering and policy, as interventions justified as “making the network safer” may shift harm across stress regimes.
We formalize this observation by defining a stress model as a triple $S=(\Omega,D,O)$, where $\Omega$ is the set of admissible perturbations, $D$ is the dynamical rule mapping perturbations to outcomes, and $O$ is the observable of interest [5, 6]. Resilience is then a functional $R(G,S)=A(\{O(D(G,\omega)):\omega\in\Omega\})$, where $A$ is an aggregation operator. This functional is undefined until $S$ and $A$ are specified. Two canonical stress models crystallize the point: (i) bond percolation, where edges are retained with probability $p$ and the observable is the giant component size, yielding the critical threshold $p_{c}$ as a fragility metric; and (ii) the Watts threshold cascade [7], where nodes activate when a fraction $\varphi$ of their neighbors are active, and the observable is the final cascade size.
Two research traditions have characterized resilience under these respective stress models largely in parallel. The percolation literature [8, 9, 10] established that scale-free networks are simultaneously robust to random failures and vulnerable to targeted attacks. The cascade literature [7, 11, 12] established that threshold-based contagion can spread globally even from small seeds. A critical question at their intersection is: when a network intervention improves resilience under one stress model, what happens under the other? This question has practical urgency. Infrastructure engineers who decommission large substations, financial regulators who break up systemically important institutions, and public health officials who target highly connected individuals are all making interventions that assume a specific failure model. If these interventions inadvertently worsen resilience under a different stress model, consequences could be severe—as demonstrated by the 2008 financial crisis, where interconnection that appeared robust to idiosyncratic defaults proved catastrophically fragile under cascade dynamics [13].
Prior work has shown qualitatively that complex networks can be “more robust and more fragile” depending on the stress model [9, 14], but stopped short of characterizing the mechanism or providing a predictive condition for when this occurs under deliberate intervention. Here we address this gap directly, studying the canonical intervention of hub removal (top 10% of nodes by degree) in Barabási–Albert networks [8] and measuring its joint effect on percolation and cascade dynamics.
@@
    A systematic study of the joint percolation-cascade phase structure under hub removal, revealing compounding vulnerability—both stress models worsen simultaneously at intermediate cascade thresholds—a phenomenon that neither stress model alone would predict.
@@
    A controlled experimental decomposition showing that hub-mediated cascade suppression is primarily dynamical (threshold-based) rather than structural (connectivity-based), demonstrated by the ${\sim}95\%/{\sim}19\%/{\sim}0.3\%$ cascade split across hub vulnerability conditions.
@@
    A closed-form analytical derivation of the post-removal cascade branching factor $z_{1}(\varphi,\alpha,m)$ under the configuration-model approximation, providing a predictive expression for cascade window expansion under targeted hub removal in BA networks.
@@ II.3 Bond Percolation Threshold
We estimate the bond percolation threshold $p_{c}$ using the $S=0.5$ method: retaining each edge independently with probability $p$ and computing the relative giant component size $S(p)=|\mathrm{GCC}|/N$. We define $p_{c}$ as the occupation probability where $S(p)=0.5$ (interpolated by bisection), sampling 200 values of $p\in[0,1]$ and averaging over 30 independent network realizations. For post-removal networks, $S(p)$ is normalized by the remaining node count $N^{\prime}=(1-\alpha)N$, ensuring comparable giant component fractions. Bootstrap 95% confidence intervals are computed with 10,00010{,}000 resamples.
This simulation-based $p_{c}$ differs from the Molloy–Reed analytical estimate $p_{c}^{\mathrm{MR}}=\langle k\rangle/(\langle k^{2}\rangle-\langle k\rangle)$ by a factor of ${\sim}3.5\times$ for BA networks, because BA networks exhibit disassortativity and finite-size effects not captured by the configuration-model assumption. We use simulation values throughout for consistency.
@@ IV.1 Three Regimes
Regime I ($\varphi<0.15$): Hub removal reduces cascades. At low thresholds, nearly all nodes are vulnerable ($k\varphi<1$ for most). Both networks are supercritical. Hub removal reduces cascade size (86.5% $\to$ 46.4% at $\varphi=0.10$) because connectivity reduction limits cascade reach.
Regime II ($0.15<\varphi<0.27$): Phase transition zone. Hub removal shifts the network from subcritical to supercritical cascade dynamics. At $\varphi=0.22$, baseline cascade size is $0.28\%$ (subcritical) versus $18.7\%$ after hub removal (supercritical), a ${\sim}67$-fold increase. This constitutes a genuine phase transition, confirmed by bimodal cascade size distributions (Fig. 2): the probability of a global cascade ($C>10\%$) rises from 0/200 before removal (95% upper bound ${\sim}1.8\%$) to 61/200 after removal ($P_{\mathrm{global}}=0.305$, Wilson 95% CI: $[0.245,0.372]$, $n=200$). Of post-removal trials, 120/200 produce cascades $1\%$ (subcritical mode) and 61/200 produce cascades $10\%$ (supercritical mode), with few intermediate values—the hallmark of a system straddling the critical point.
Figure 2: Cascade size distributions before (left) and after (right) hub removal at $\varphi=0.22$, BA(N=2, [^0] N=2{,}000, $m=2$, 200 trials). Before removal, cascades are unimodal near zero. After removal, the distribution becomes bimodal, with modes at $1\%$ and $10\%$—the hallmark of a phase transition straddling the critical point.
Regime III ($\varphi>0.27$): Both subcritical. Nearly no nodes are vulnerable. Hub removal modestly worsens cascades (relative increases) but absolute sizes remain $0.5\%$.
@@ IV.2 Cascade Response Across $\varphi$
Figure 3 plots the mean cascade size before and after hub removal as a function of $\varphi$ (log scale). Regime I corresponds to low thresholds where both networks are globally supercritical and hub removal reduces cascade reach by deleting connectivity. Regime II is a narrow transition zone where hub removal induces large global cascades at previously safe operating points. The most striking effect occurs near $\varphi\in[0.22,0.25]$, where the baseline cascade size is $0.3\%$ but the post-removal mean is ${\sim}20\%$. Regime III corresponds to high thresholds where both networks are subcritical and cascade sizes remain $0.5\%$.
@@ IV.3 Compounding Vulnerability
Hub removal produces compounding vulnerability: both percolation and cascade metrics worsen simultaneously. The percolation threshold increases by $168\%$ ($p_{c}$: $0.337\to 0.903$), meaning the remaining network requires $2.68\times$ more edge retention for connectivity. Simultaneously, the cascade window expands, enabling global cascades at previously safe operating points. This compounding—rather than a trade-off—is the central finding: hub removal does not exchange one type of resilience for another but degrades both.
As expected from model independence, $\Delta F_{p}\approx+168\%$ is independent of cascade threshold $\varphi$, since percolation depends only on the static degree sequence. We report $\Delta F_{p}$ once as a baseline structural effect of hub removal.
@@ IV.4 Topology Dependence and the WS Cliff-Edge
The cascade window expansion is mediated by degree heterogeneity. In Watts–Strogatz networks, the firewall population exhibits a dramatic cliff-edge transition absent in BA networks. At $\varphi=0.25$, only 15% of WS nodes are stable ($k\geq 5$). At $\varphi=0.30$, the threshold drops to $k\geq 4$, and since most WS nodes have $k=4$, the stable fraction jumps to 84% in a single $\varphi$ step. This means WS networks are safe across a wide $\varphi$ range, then suddenly vulnerable at a single threshold—a catastrophic cliff-edge.
@@ 4 ${\sim}17$ 7/10 10/10 10/10 10/10
The transition is sharp: for $m=1$ ($\kappa\approx 8$), hub removal rarely worsens cascades and often improves them. For $m\geq 2$ ($\kappa\geq 11$), compounding vulnerability emerges robustly across the $(m,\alpha)$ grid. This confirms the $\kappa>{\sim}10$ threshold and demonstrates that the phenomenon is not an artifact of a single parameter choice.
@@ IV.6 Finite-Size Scaling
The percolation fragility increase is remarkably stable: $\Delta F_{p}=+168\%$ at N=2, [^0] N=2{,}000, $167\%$ at N=5, [^0] N=5{,}000, $166\%$ at N=10, [^0] N=10{,}000. The cascade phenomenon persists across N=500N=50010,00010{,}000, with the transition sharpening as $N$ increases (see repository datasets and scripts in Appendix C).
@@ V.1 Stable-Node Condition
Proposition 1 (Stable-node condition). A node $v$ with degree $k_{v}$ blocks single-neighbor cascade propagation at threshold $\varphi$ if and only if $k_{v}>1/\varphi$.
This follows directly from the Watts activation rule: node $v$ requires $\lceil k_{v}\cdot\varphi\rceil$ active neighbors, which exceeds 1 when $k_{v}>1/\varphi$. The fraction of stable nodes is $f_{\mathrm{stable}}(\varphi)=P(k>1/\varphi)$. For BA networks at $\varphi=0.22$: $f_{\mathrm{stable}}\approx 0.20$.
@@ V.2 Lower Bound on the Cascade Transition
Proposition 2 (Necessary condition for cascade window expansion). For BA($m$) networks with hub removal fraction $\alpha$, the cascade transition threshold satisfies $\varphi^{*}\geq\sqrt{\alpha}/m$.
Proof. The minimum degree among removed hubs scales as $k_{\min}^{\mathrm{hub}}\approx m/\sqrt{\alpha}$ (from BA degree ranking $k_{r}\approx m\sqrt{N/r}$ evaluated at $r=\alpha N$). For cascade window expansion, removed hubs must be stable nodes: $k_{\min}^{\mathrm{hub}}>1/\varphi$, yielding $\varphi>\sqrt{\alpha}/m$. ∎
For $m=2$, $\alpha=0.10$: $\varphi^{*}_{\mathrm{lower}}=\sqrt{0.10}/2\approx 0.158$. The observed transition at $\varphi^{*}\approx 0.17$ is within 8%—this bound is network-size independent and computable from construction parameters alone.

### 07 Mar 2026 v2

# Compounding Vulnerability: Hub Removal Can Trigger Cascade Phase Transitions While Degrading Percolation Robustness in Barabási–Albert Networks

@@ Abstract
Targeted hub removal is known to weaken connectivity in heterogeneous networks. We show that in Barabási–Albert networks the same intervention can also shift Watts threshold dynamics across the cascade critical point. For BA networks with N=2, [^0] N=2{,}000 and $m=2$, removing the top 10% of nodes by degree raises the bond-percolation threshold from $p_{c}=0.174$ to $0.776$ and, at $\varphi=0.22$, increases mean cascade size from $0.86\%$ (95% CI 0.43–1.30) to $23.1\%$ (21.3–24.9). A controlled hub-vulnerability experiment on fixed topology shows that most of this cascade effect is dynamical: lowering hub activation thresholds produces much larger cascades even without deleting nodes, while deletion partly offsets the increase by removing edges. Using a configuration-model approximation, we derive the post-removal branching factor $z_{1}$ and identify a window in which the original network is subcritical but the hub-removed network is supercritical. The effect persists across system sizes and is not seen in matched ER or WS controls. These results identify a regime in which hub removal simultaneously worsens connectivity and cascade exposure in BA networks.
@@ I Introduction
Network resilience—the ability of a system to maintain function under stress—is a central problem of complexity science [1, 2, 3, 27, 28]. Resilience is not a property of a graph $G$ alone, but of the pair $(G,S)$ consisting of a network and a specified stress model [4]. Two canonical stress models illustrate this [5, 6]: (i) bond percolation, where edges are retained with probability $p$ and the observable is the giant component size, yielding the critical threshold $p_{c}$ as a fragility metric; and (ii) the Watts threshold cascade [8], where nodes activate when a fraction $\varphi$ of their neighbors are active, and the observable is the final cascade size.
Two research traditions have characterized resilience under these respective stress models largely in parallel. The percolation literature [9, 10, 39, 11, 30] established that scale-free networks are simultaneously robust to random failures and vulnerable to targeted attacks. The cascade literature [8, 12, 13, 31, 32] established that threshold-based contagion can spread globally even from small seeds, with recent extensions to temporal networks and heterogeneous thresholds. A critical question at their intersection is: when a network intervention improves resilience under one stress model, what happens under the other? This question has practical urgency. Infrastructure engineers who decommission large substations, financial regulators who break up systemically important institutions, and public health officials who target highly connected individuals are all making interventions that assume a specific failure model. If these interventions inadvertently worsen resilience under a different stress model, consequences could be severe—as demonstrated by the 2008 financial crisis, where interconnection that appeared robust to idiosyncratic defaults proved catastrophically fragile under cascade dynamics [14].
Prior work has shown qualitatively that complex networks can be “more robust and more fragile” depending on the stress model [10, 15], and recent studies have explored cascading failures across interdependent networks [28, 35] and scaling laws of failure dynamics [29], but stopped short of characterizing the mechanism or providing a predictive condition for when compounding vulnerability occurs under deliberate intervention on a single network. Here we address this gap directly, studying the canonical intervention of hub removal (top 10% of nodes by degree) in Barabási–Albert networks [9] and measuring its joint effect on percolation and cascade dynamics.
@@
    A systematic study of the joint percolation-cascade phase structure under hub removal in BA networks, demonstrating that both stress models can worsen simultaneously at intermediate cascade thresholds—a phenomenon that neither stress model alone would predict.
@@
    A controlled experimental decomposition showing that hub-mediated cascade suppression is primarily dynamical (threshold-based) rather than structural (connectivity-based), demonstrated by the ${\sim}95\%/{\sim}19\%/{\sim}0.3\%$ cascade split across hub vulnerability conditions on the same network topology.
@@
    A closed-form analytical derivation of the post-removal cascade branching factor $z_{1}(\varphi,\alpha,m)$ under the configuration-model approximation, providing a predictive expression for cascade window expansion under targeted hub removal.
@@ II.3 Bond Percolation Threshold
We estimate the bond percolation threshold $p_{c}$ using the susceptibility peak method [7]: for each occupation probability $p$, edges are retained independently with probability $p$ and we compute the mean finite-cluster size $\chi(p)=\sum_{s}s^{2}n_{s}/\sum_{s}s\cdot n_{s}$ (excluding the giant component), where $n_{s}$ is the number of clusters of size $s$. The percolation threshold is defined as $p_{c}=\arg\max_{p}\chi(p)$, the standard estimator in computational percolation studies. We sample 200 values of $p\in[0,1]$ with 50 bond realizations per $p$-value, averaged over 30 independent network realizations. Bootstrap 95% confidence intervals are computed with 10,00010{,}000 resamples.
As a robustness check, we also compute $p_{c}$ using the $S=0.5$ crossing method (where the relative giant component size $S(p)=|\mathrm{GCC}|/N$ reaches $0.5$) and the Molloy–Reed analytical estimate $p_{c}^{\mathrm{MR}}=\langle k\rangle/(\langle k^{2}\rangle-\langle k\rangle)$. All three methods yield consistent directional results ($\Delta F_{p}\gg 0$), though quantitative values differ due to BA disassortativity and finite-size effects (see Appendix A for comparison). We report susceptibility-peak values throughout.
@@ IV.1 Three Regimes
Regime I ($\varphi\leq 0.20$): Both supercritical; hub removal reduces cascades. At low thresholds, most nodes satisfy the vulnerability condition $\lfloor k\varphi\rfloor\geq 1$. Both networks sustain large cascades. Hub removal reduces cascade size (90.0% $\to$ 42.6% at $\varphi=0.10$; 33.3% $\to$ 40.1% at $\varphi=0.20$) because connectivity reduction limits cascade reach. The $z_{1}$ table (Table 6) confirms: at $\varphi=0.18$$0.20$, both pre- and post-removal $z_{1}>1$.
Regime II ($0.20<\varphi\leq 0.25$): Phase transition zone. Hub removal opens a cascade window where the pre-removal network is subcritical but the post-removal network is supercritical. The boundary at $\varphi\approx 0.20$ is set by the integer threshold $\lfloor 1/\varphi\rfloor$ crossing from 5 to 4 (Table 6). At $\varphi=0.22$, mean cascade size increases from $0.86\%$ [95% CI: 0.43–1.30] to $23.1\%$ [95% CI: 21.3–24.9] ($n_{\mathrm{eff}}=50$ independent networks, 20 seeds each; total n=1,000n=1{,}000; within-network seeds are correlated, so network-level effective sample size is $n_{\mathrm{eff}}\approx 50$). This constitutes a genuine phase transition, confirmed by bimodal cascade size distributions (Fig. 2): the probability of a global cascade ($C>10\%$) rises from 0/1,000 before removal to 377/1,000 after removal ($P_{\mathrm{global}}=0.377$, Wilson 95% CI: $[0.348,0.406]$). Post-removal trials split between a subcritical mode ($C<1\%$) and a supercritical mode ($C>10\%$), with few intermediate values—the hallmark of a system straddling the critical point. The hub vulnerability experiment (Table 2, $n=200$: 10 networks $\times$ 20 seeds) yields directionally consistent results ($0.28\%\to 18.7\%$); the lower baseline (0.28% vs. 0.86%) reflects higher variance at small effective sample ($n_{\mathrm{eff}}=10$ networks) and is order-of-magnitude consistent with the main experiment.
Figure 2: Cascade size distribution after hub removal at $\varphi=0.22$, BA(N=2, [^0] N=2{,}000, $m=2$, 200 trials). The main panel shows the post-removal distribution, which is bimodal with modes at $1\%$ (subcritical) and $10\%$ (global cascade). The inset shows the pre-removal distribution, which is unimodal near zero—all cascades subcritical. The transition from unimodal to bimodal is the signature of a phase transition straddling the critical point.
Regime III ($\varphi>0.25$): Both subcritical. At $\varphi=0.26$, even post-removal $z_{1}<1$ (Table 6). Hub removal modestly worsens cascades (relative increases) but absolute sizes remain $0.5\%$.
@@ IV.2 Cascade Response Across $\varphi$
Figure 3 plots the mean cascade size before and after hub removal as a function of $\varphi$ (log scale). Regime I ($\varphi\leq 0.20$) corresponds to thresholds where both networks are supercritical and hub removal reduces cascade reach by deleting connectivity. Regime II ($0.20<\varphi\leq 0.25$) is a narrow transition zone where hub removal induces large global cascades at previously subcritical operating points. The most striking effect occurs near $\varphi\in[0.21,0.25]$, where the baseline cascade size is $1\%$ but the post-removal mean exceeds $20\%$. Regime III ($\varphi>0.25$) corresponds to high thresholds where both networks are subcritical and cascade sizes remain $0.5\%$.
@@ IV.3 Compounding Vulnerability
Hub removal produces compounding vulnerability: both percolation and cascade metrics worsen simultaneously. The percolation threshold increases substantially, meaning the remaining network requires far more edge retention for connectivity. Simultaneously, the cascade window expands, enabling global cascades at previously safe operating points. This compounding—rather than a trade-off—is the central finding: hub removal does not exchange one type of resilience for another but degrades both.
As expected from model independence, $\Delta F_{p}$ is independent of cascade threshold $\varphi$, since percolation depends only on the static degree sequence. We report two standard estimators for transparency:
  • Susceptibility peak [7]: $\Delta F_{p}=+347\%$ ($p_{c}$: $0.174\pm 0.009\to 0.776\pm 0.036$, 30 networks $\times$ 200 $p$-values);
  • $S(p)=0.5$ crossing: $\Delta F_{p}=+170\%$ ($p_{c}$: $0.331\to 0.895$).
The factor-of-two difference between estimators arises because the susceptibility peak locates the onset of the phase transition (where fluctuations diverge), while $S=0.5$ locates the midpoint (where half the network is connected). Both are standard; neither is “correct” in isolation. The qualitative conclusion—massive percolation degradation—is robust to estimator choice. We report susceptibility-peak values as primary throughout, consistent with [7].
@@ IV.4 Topology Dependence and the WS Cliff-Edge
The cascade window expansion requires sufficient degree heterogeneity. In homogeneous Watts–Strogatz networks, the stable-node fraction undergoes a cliff-edge transition (not observed in BA): at $\varphi=0.25$, only 15% of WS nodes are stable ($k\geq 5$), but at $\varphi=0.30$, the threshold drops to $k\geq 4$ and the stable fraction jumps to 84% in a single step—a consequence of the narrow degree distribution centered on $k=4$. This cliff-edge prevents the gradual firewall degradation that drives the BA phase transition.
@@ 4 ${\sim}17$ 7/10 10/10 10/10 10/10
The transition is qualitatively sharp: for $m=1$ ($\kappa\approx 8$), hub removal rarely worsens cascades and often improves them. For $m\geq 2$ ($\kappa\geq 11$), compounding vulnerability emerges across the $(m,\alpha)$ grid. With only 10 realizations per cell, the exact fractions should be interpreted as indicators of the trend rather than precise probability estimates; the qualitative pattern ($m\geq 2$ robustly worsens, $m=1$ does not) is the key finding. This confirms the $\kappa>{\sim}10$ threshold.
@@ IV.6 Finite-Size Scaling
The percolation fragility increase is stable—and in fact increasing—across network sizes under the susceptibility-peak method ($8$ networks per size):
Table 5: Finite-size scaling of percolation fragility (susceptibility-peak method, 8 networks per $N$).
NN $p_{c}^{\mathrm{pre}}$ $p_{c}^{\mathrm{post}}$ $\Delta F_{p}$
2,000 $0.171\pm 0.017$ $0.773\pm 0.038$ $352\%$
5,000 $0.157\pm 0.009$ $0.759\pm 0.027$ $383\%$
10,000 $0.146\pm 0.008$ $0.749\pm 0.017$ $412\%$
The N=2, [^0] N=2{,}000 estimate ($\Delta F_{p}=+352\%$) reported throughout is therefore conservative: larger networks show even stronger percolation degradation, consistent with finite-size corrections decaying as $O(1/\ln N)$. The narrowing standard deviations confirm convergence. The cascade phenomenon persists across N=500N=50010,00010{,}000, with the transition sharpening as $N$ increases (see repository datasets and scripts in Appendix E).
@@ V.1 Stable-Node Condition
Definition 1 (Stable node). A node $v$ with degree $k_{v}$ is stable (cascade-resistant) at threshold $\varphi$ when $k_{v}>1/\varphi$, since it then requires $\lceil k_{v}\cdot\varphi\rceil\geq 2$ active neighbors to activate—i.e., a single active neighbor cannot trigger it. The fraction of stable nodes is $f_{\mathrm{stable}}(\varphi)=P(k>1/\varphi)$. For BA networks at $\varphi=0.22$: $f_{\mathrm{stable}}\approx 0.20$.
@@ V.2 Lower Bound on the Cascade Transition
Observation 1 (Necessary condition for cascade window expansion). For BA($m$) networks with hub removal fraction $\alpha$, the lower onset of the cascade window satisfies $\varphi_{\mathrm{onset}}\geq\sqrt{\alpha}/m$.
Derivation. The minimum degree among removed hubs scales as $k_{\min}^{\mathrm{hub}}\approx m/\sqrt{\alpha}$ (from the BA degree ranking approximation $k_{r}\approx m\sqrt{N/r}$ evaluated at $r=\alpha N$; this is a thermodynamic-limit scaling argument with $O(1/\ln N)$ finite-size corrections). For cascade window expansion, removed hubs must be stable nodes: $k_{\min}^{\mathrm{hub}}>1/\varphi$, yielding $\varphi>\sqrt{\alpha}/m$. For $m=2,\alpha=0.10$: $\varphi_{\mathrm{onset}}\geq 0.158$, consistent with the observed transition at $\varphi\approx 0.17$.
For $m=2$, $\alpha=0.10$: $\varphi_{\mathrm{onset}}=\sqrt{0.10}/2\approx 0.158$. The observed transition at $\varphi\approx 0.17$ is within 8%—this bound is network-size independent and computable from construction parameters alone.